10. Omar has Cobb Douglas preferences for Juice (X1) and Soda (X2). His MRS at his optimal consumption is 2. If the price of juice is p1= $6. What is the price of Soda, p2?
Answer #1

At the optimal consumption,

MRS = price ratio ( Price of juice / Price of Soda)

Put values, we get

MRS = p1/ p2

2 = 6/ p2

p2 = 6/2 = 3

So price of Soda, P2 = $3.